OpenStudy (anonymous):

the summ of a certain number and a second number is -42. the first number minus the second is 52. Find the two numbers. HOW DO I SOLVE QUESTIONS LIKE THIS?? AND WHATS THE ANSWER?????

OpenStudy (anonymous):

a + b = -42 a - b = 52 add the two equations: a + b + a - b = -42 + 52 so 2a = 10 so a = 5

hero (hero):

x + y = -42 x - y = 52 x = 5 y = -47

hero (hero):

Sorry I kind of skipped some steps, but if you do the algebra, you should be about to get it.

OpenStudy (anonymous):

therefore b = -47